A chemist requires 0.250 mol Na2CO3 for a reaction. How many grams does this correspond to?
1
Expert's answer
2019-07-22T07:22:53-0400

M (Na2CO3) = 2 × 23 + 12 + 3 × 16 = 106 g mol-1

m (Na2CO3) = 0.250 × 106 = 26.5 g